OpenStudy (anonymous):

A pendulum with a period of 1.8 s on Earth is moved to another planet. On the new planet, its period is 2.5 s. What can you conclude about this new planet?

OpenStudy (hoothoot):

The force of gravity on the new planet is less than it is on Earth.

OpenStudy (hoothoot):

that is your correct answer

OpenStudy (irishboy123):

\[T = 2\pi \sqrt{\frac{L}{g}}\]